lim [n 2 -n+1/n 2 -n-1] n(n-1) n_# # is infinity

 
lim [n2-n+1/n2-n-1]n(n-1)
n_#
# is infinity

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
1^(infinite form)
lim n_infinite[n2-n+1/n2-n-1]^(n(n-1))

=lim n_infinite(1+2/n2-n-1)^n(n-1)
=e^(lim n_infinite 2n(n-1)/n2-n-1)
=e^(lim n_infinite 2-2/n)/1-1/n-1/n2)
=e^2

Sher Mohammad
Faculty askiitians